# On the Convergence of Multi-objective Optimization under Generalized Smoothness

### v1

@@ Abstract
Multi-objective optimization (MOO) is receiving more attention in various fields such as multi-task learning. Recent works provide some effective algorithms with theoretical analysis but they are limited by the standard $L$-smooth or bounded-gradient assumptions, which are typically unsatisfactory for neural networks, such as recurrent neural networks (RNNs) and transformers. In this paper, we study a more general and realistic class of $\ell$-smooth loss functions, where $\ell$ is a general non-decreasing function of gradient norm. We develop two novel single-loop algorithms for $\ell$-smooth MOO problems, Generalized Smooth Multi-objective Gradient descent (GSMGrad) and its stochastic variant, Stochastic Generalized Smooth Multi-objective Gradient descent (SGSMGrad), which approximate the conflict-avoidant (CA) direction that maximizes the minimum improvement among objectives. We provide a comprehensive convergence analysis of both algorithms and show that they converge to an $\epsilon$-accurate Pareto stationary point with a guaranteed $\epsilon$-level average CA distance (i.e., the gap between the updating direction and the CA direction) over all iterations, where totally $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples are needed for deterministic and stochastic settings, respectively. Our algorithms can also guarantee a tighter $\epsilon$-level CA distance in each iteration using more samples. Moreover, we propose a practical variant of GSMGrad named GSMGrad-FA using only constant-level time and space, while achieving the same performance guarantee as GSMGrad. Our experiments validate our theory and demonstrate the effectiveness of the proposed methods.
@@ 1.1 Our Contributions
New single-loop algorithms. Both GSMGrad and SGSMGrad are easy to implement by updating the weights $w$ of objectives and model parameters $x$ simultaneously via a a single-loop structure. A warm-start initialization sub-procedure is also introduced at the beginning of both algorithms to ensure a sufficiently small iteration-wise CA distance under the single-loop structure. We also propose a computation- and memory-efficient variant of GSMGrad named GSMGrad-FA by updating the objective weights $w$ using only forward passes of $F(\cdot)$ rather than gradient $\nabla F$, which effectively reduces $O(K)$ time and space to $O(1)$ without hurting the performance guarantee.
Convergence analysis and optimal complexity. We provide a comprehensive analysis of our proposed algorithms under the generalized $\ell$-smooth condition in both deterministic and stochastic settings. To achieve an $\epsilon$-accurate Pareto stationary point and an $\epsilon$-level average CA distance, we show that GSMGrad and SGSMGrad require $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples in the deterministic and stochastic settings, respectively. Furthermore, to achieve a more aggressive $\epsilon$-level iteration-wise CA distance, GSMGrad and SGSMGrad require an increased number of samples, on the order of $\mathcal{O}(\epsilon^{-11})$ and $\mathcal{O}(\epsilon^{-17})$ respectively, in both deterministic and stochastic scenarios due to smaller step sizes and mini-batch data sampling. Moreover, we show that GSMGrad-FA achieves the same performance guarantee as GSMGrad.
@@ 1.2 Related Works
Gradient-based multi-objective optimization. A variety of gradient manipulation techniques have emerged for simultaneous learning of multiple tasks. One prevalent category of methods adjusts the weights of various objectives according to factors such as uncertainty [20], gradient norm [7], and training complexity [17]. Methods based on MOO have garnered increased attention due to their systematic designs, enhanced training stability and model-agnostic nature. For instance, [30] framed Multi-Task Learning (MTL) as a MOO problem and introduced an optimization method akin to MGDA [12]. Afterward, many MGDA-based methods have been proposed to mitigate gradient conflict with promising empirical performance. Among them, PCGrad [34] avoids conflict by projecting the gradient of each task on the norm plane of other tasks. GradDrop [8] randomly drops out conflicted gradients. CAGrad [24] adds a constraint on the update direction to be close to the average gradient. NashMTL [27] and FairGrad [3] formulated MTL as a bargaining game and a resource allocation problem, respectively. Theoretically, [13] proposed a provably convergent stochastic MOO method named MoCo based on an auxiliary tracking variable for gradient approximation. [6] characterized the trade-off among optimization, generalization, and conflict avoidance in MOO. [32] proposed a stochastic MOO method named SDMGrad with a preference-oriented regularizer, and analyzed its convergence. However, all these works rely on the $L$-smoothness and bounded-gradient assumptions. This paper focuses on the MOO problems with generalized $\ell$-smooth objectives.
@@ 2.3 Multiple-gradient descent algorithm (MGDA) and its stochastic variants
$d_{s}^{*}=\nabla F(x;s)w_{s}^{*};\;\;s.t.\;w_{s}^{*}\in\arg\min_{w\in\mathcal{W}}\frac{1}{2}\|\nabla F(x;s)w\|^{2},$ (5)
@@ 3.3 Fast approximation via Taylor expansion
${F}(x_{t})-{F}(x_{t+1})=\nabla F(x_{t})^{\top}(x_{t}-x_{t+1})-R(x_{t})=\alpha\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}-R(x_{t}),$ (7)
@@ 4.1 Deterministic setting
Let Assumptions 1 and 2 hold. Set $\beta\sim\mathcal{O}(\frac{1}{M^{2}}),\alpha\sim\mathcal{O}(\frac{1}{M^{2}}+\frac{1}{M\ell(M+1)}),T\geq\max\left(\Theta\left(\frac{1}{\alpha\epsilon^{2}}\right),\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\right),$ and $\rho\sim\mathcal{O}(\epsilon^{2})$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
@@ Appendix D Detailed Proofs for Iteration-wise CA Distance
$+\beta\|\big{(}\nabla F(x_{t})^{\top}\nabla F(x_{t})-\nabla F^{\top}(x_{t+1})\nabla F(x_{t+1})\big{)}w_{Q,\rho}(x_{t+1})\|.$ (28)
@@ D.4 Formal Version of Its Proof of Theorem 6
$+\left(1+\frac{1}{a_{2}}+\frac{1}{a_{3}}\right)\alpha^{2}L_{w}^{2}\left(2M^{2}+\frac{4\sigma^{2}}{n_{s}}\right)$

### v2

@@ Abstract
Multi-objective optimization (MOO) is receiving more attention in various fields such as multi-task learning. Recent works provide some effective algorithms with theoretical analysis but they are limited by the standard $L$-smooth or bounded-gradient assumptions, which are typically unsatisfactory for neural networks, such as recurrent neural networks (RNNs) and transformers. In this paper, we study a more general and realistic class of $\ell$-smooth loss functions, where $\ell$ is a general non-decreasing function of gradient norm. We develop two novel single-loop algorithms for $\ell$-smooth MOO problems, Generalized Smooth Multi-objective Gradient descent (GSMGrad) and its stochastic variant, Stochastic Generalized Smooth Multi-objective Gradient descent (SGSMGrad), which approximate the conflict-avoidant (CA) direction that maximizes the minimum improvement among objectives. We provide a comprehensive convergence analysis of both algorithms and show that they converge to an $\epsilon$-accurate Pareto stationary point with a guaranteed $\epsilon$-level average CA distance (i.e., the gap between the updating direction and the CA direction) over all iterations, where totally $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples are needed for deterministic and stochastic settings, respectively. Our algorithms can also guarantee a tighter $\epsilon$-level CA distance in each iteration using more samples. Moreover, we propose an efficient variant of GSMGrad named GSMGrad-FA using only constant-level time and space, while achieving the same performance guarantee as GSMGrad. Our experiments validate our theory and demonstrate the effectiveness of the proposed methods.
@@ 1.1 Our Contributions
New single-loop algorithms. Both GSMGrad and SGSMGrad are easy to implement by updating the weights $w$ of objectives and model parameters $x$ simultaneously via a single-loop structure. A warm-start initialization sub-procedure is also introduced at the beginning of both algorithms to ensure a sufficiently small iteration-wise CA distance under the single-loop structure. We also propose a computation- and memory-efficient variant of GSMGrad named GSMGrad-FA by updating the objective weights $w$ using only forward passes of $F(\cdot)$ rather than gradient $\nabla F$, which effectively reduces $O(K)$ time and space to $O(1)$ without hurting the performance guarantee.
Convergence analysis and optimal complexity. We provide a comprehensive analysis of our proposed algorithms under the generalized $\ell$-smooth condition in both deterministic and stochastic settings. To achieve an $\epsilon$-accurate Pareto stationary point and an $\epsilon$-level average CA distance, we show that GSMGrad and SGSMGrad require $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ samples in the deterministic and stochastic settings, respectively. Furthermore, to achieve a more aggressive $\epsilon$-level iteration-wise CA distance, GSMGrad and SGSMGrad require an increased number of samples, on the order of $\mathcal{O}(\epsilon^{-11})$ and $\mathcal{O}(\epsilon^{-17})$ respectively, in both deterministic and stochastic scenarios due to smaller step sizes and mini-batch data sampling. Typically, achieving an $\epsilon$-level iteration-wise CA distance results in much higher sample complexity, such as $\mathcal{O}(\epsilon^{-24})$ in [13], $\mathcal{O}(\epsilon^{-16})$ in [6] and $\mathcal{O}(\epsilon^{-12})$ in [32] for non-convex stochastic setting. Moreover, we show that GSMGrad-FA achieves the same performance guarantee as GSMGrad.
@@ 1.2 Related Works
Gradient-based multi-objective optimization. A variety of gradient manipulation techniques have emerged for simultaneous learning of multiple tasks. One prevalent category of methods adjusts the weights of various objectives according to factors such as uncertainty [20], gradient norm [7], and training complexity [17]. Methods based on MOO have garnered increased attention due to their systematic designs, enhanced training stability and model-agnostic nature. For instance, [30] framed Multi-Task Learning (MTL) as a MOO problem and introduced an optimization method akin to MGDA [12]. Afterward, many MGDA-based methods have been proposed to mitigate gradient conflict with promising empirical performance. Among them, PCGrad [34] avoids conflict by projecting the gradient of each task on the norm plane of other tasks. GradDrop [8] randomly drops out conflicted gradients. CAGrad [24] adds a constraint on the update direction to be close to the average gradient. NashMTL [27] and FairGrad [3] formulated MTL as a bargaining game and a resource allocation problem, respectively. Theoretically, [13] proposed a provably convergent stochastic MOO method named MoCo based on an auxiliary tracking variable for gradient approximation. [6] characterized the trade-off among optimization, generalization, and conflict avoidance in MOO. [32] proposed a stochastic MOO method named SDMGrad with a preference-oriented regularizer, and analyzed its convergence. However, all these works rely on the $L$-smoothness and bounded-gradient assumptions. The details can be founded in Table 1. This paper focuses on the MOO problems with generalized $\ell$-smooth objectives.
Method Smoothness 11footnotemark: 1 Assumption22footnotemark: 2 Sample Complexity
SMG [25] (LS) (BG) N/A33footnotemark: 3
CR-MOGM[36] (LS) (BF), (BG) $\mathcal{O}(\epsilon^{-4})$
MoCo[13] (LS) (BF), (BG) $\mathcal{O}(\epsilon^{-4})$
MoDo[6] (LS) (BG) $\mathcal{O}(\epsilon^{-4})$
SDMGrad[32] (LS) (BG) $\mathcal{O}(\epsilon^{-4})$
SGSMGrad (this paper) (GS) N/A $\mathcal{O}(\epsilon^{-4})$
Table 1: Comparison for existing stochastic methods for MOO problems to obtain an $\epsilon$–accurate Pareto stationary point. Explanation on the upper footmarks: $1:$ (LS) indicates that the objectives are standard $L$-smooth while (GS) the objectives are generalized $\ell$-smooth defined in Definition 1. $2:$ (BF) shows that the bounded function value assumption is required and (BG) indicates an additional bounded assumption is required. $3:$ the method in [25] only works for convex objectives and it is unable to obtain an $\epsilon$–accurate Pareto stationary point for non-convex objectives studied in this paper.
@@ 2.3 Multiple-gradient descent algorithm (MGDA) and its stochastic variants
$d_{s}^{*}=\nabla F(x;s)w_{s}^{*};\;\;s.t.\;w_{s}^{*}\in\arg\min_{w\in\mathcal{W}}\frac{1}{2}\|\nabla F(x;s)w\|^{2},$
@@ 3.3 Fast approximation via Taylor expansion
${F}(x_{t})-{F}(x_{t+1})=\nabla F(x_{t})^{\top}(x_{t}-x_{t+1})-R(x_{t})=\alpha\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}-R(x_{t}),$
@@ 4.1 Deterministic setting
Let Assumptions 1 and 2 hold. Set $\beta\sim\mathcal{O}(\frac{1}{M^{2}}),\alpha\sim\mathcal{O}(\frac{1}{M^{2}}+\frac{1}{M\ell(M+1)}),T\geq\max\Big{(}\Theta\big{(}\frac{1}{\alpha\epsilon^{2}}\big{)},\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\Big{)}$ and $\rho\sim\mathcal{O}(\epsilon^{2})$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
@@ Appendix D Detailed Proofs for Iteration-wise CA Distance
$+\beta\|\big{(}\nabla F(x_{t})^{\top}\nabla F(x_{t})-\nabla F^{\top}(x_{t+1})\nabla F(x_{t+1})\big{)}w_{Q,\rho}(x_{t+1})\|.$
@@ D.4 Formal Version of Its Proof of Theorem 6
$+\left(1+\frac{1}{a_{2}}+\frac{1}{a_{3}}\right)\alpha^{2}L_{w}^{2}\left(2M^{2}+\frac{4\sigma^{2}}{n_{s}}\right),$ (46)